85. A parallel-plate capacitor with plate area 3.0 cm2 and air- gap separation 0.50 mm is connected to a 12-V battery, and fully charged. The battery is then disconnected. (a) What is the charge on the capacitor? (b) The plates are now pulled to a separation of 0.75 mm. What is the charge on the capacitor now? (c) What is the potential difference between the plates now? (d) How much work was required to pull the plates to their new separation?
media%2F334%2F334ba347-d329-4764-a588-ce
Answer #1

Given,

A = 3 cm^2 = 3 x 10^-4 m^2 ; d = 0.5 mm = 0.5 x 10^-3 m ; V = 12 V

a)We know that

C = e0 A/d

C = 8.85 x 10^-12 x 3 x 10^-4/0.5 x 10^-3 = 5.31 x 10^-12 F

Q = CV = 5.31 x 10^-12 x 12 = 63.72 pF

Hence, Q = 63.72 x 10^-12 F = 63.72 pF

b)The charge will remains same, as conservation of charge applies here and there is no path for the discharge available

Q = 63.72 pC

c)C = 8.85 x 10^-12 x 3 x 10^-4/0.75 x 10^-3 = 3.54 x 10^-12 F

Q = CV => V = Q/C

V = 63.72/3.54 = 18 V

Hence. V = 18 V

d)The work required will be:

W = Uf - Ui

W = 1/2 Q (V2 - V1)

W = 0.5 x 63.72 x 10^-12 x (18 - 12) = 1.91 x 10^-10 J

Hence, W = 1.91 x 10^-10 J

[kindly check and rate, let me know for any query. Pls put one question at a time as per the guidelines. Thanks and all the best]